# SOAC: Spatio-Temporal Overlap-Aware Multi-Sensor Calibration using Neural Radiance Fields

### 27 Nov 2023 v1

@@ 3.6 Optimization Details
Overall, the training process can be summarized as follows: during each training step, a mini-batch of rays is first used in the scene representation training step (Section 1(a)). Rays of each camera train their respective NeRF and fill the respective visibility grids (Section 3.5). The LiDAR rays train all the NeRFs and are used to optimize the LiDAR calibration parameters after being filtered by the visibility grids. In a subsequent step, the same mini-batch is passed to the extrinsic and temporal optimization. Rays are filtered through the visibility grids before being fed to the NeRFs as explained in section 3.3. Calibration losses (equation 8) are computed and the gradient is backpropagated to optimize the calibration parameters. Once this is done, we continue the training with the next mini-batch.
@@
Figure 5: Results for SOAC and MOISST as box plots with log scale on KITTI-360 and Nuscenes sequences. The red lines show the initial error (Best viewed in color).
@@ 4.1 Setup
For our experiments, we use three public driving scene datasets: KITTI-360 [18], nuScenes [3] and Pandaset [42]. For KITTI-360, we consider the two front cameras, the two side cameras and the Velodyne LiDAR. For nuScenes and Pandaset, we consider the front camera, the two front diagonal cameras, and the LiDAR. We assign the front-left camera of KITTI-360, and the front camera of nuScenes and Pandaset to be the reference sensor. More details on selected sequences and dataset parameters are provided in the supplementary materials.
@@
As a comparison baseline we use MOISST, as it is also a method that aims to solve the targetless, multi-modal and spatiotemporal calibration problem. We implemented the method as described by the paper [12] with slight changes (more details in supplementary materials). For SOAC, we use the same architecture as MOISST for the NeRF model and apply the same supervision and regularization losses. More details on the hyperparameters used are given in the supplementary materials. For the LiDAR/Camera calibration task, we compare with LCCNet [20] by using the code and the pretrained weights from the official repository [^1] https://github.com/IIPCVLAB/LCCNet.
@@ 4.2 Results
We run the benchmarks on 4 KITTI-360 sequences and 3 nuScenes sequences. For SOAC, we downscale by a factor of 4 the image dimensions of KITTI-360 and by a factor of 6 for nuScenes. For MOISST, we do not downscale the KITTI-360 images and downscale by a factor of 2 nuScenes images as this method performs better with high-resolution images. Each test is run with an initial noise of 50 cm translation error and 5° rotation error on each axis, as well as 100 ms of time offset. We use 10 different seeds to randomly add positive or negative signs for each added noise and compute the error statistics over these 10 runs. We provide the box plots [^2] The boxes show the first quartile $Q_{1}$, median, third quartile $Q_{3}$. The whiskers use 1.5 IQR (Interquartile range) above and below the box and stop at a value within the results of the results in Figure 5. We obtain better calibration results on KITTI-360 an overall error (average over median for each sensor: rotation/translation/time offset) for SOAC: 0.21°/5.24 cm/3.95 ms compared to MOISST: 0.21°/10.02 cm/4.19 ms. On nuScenes, the difference is wider as we have for SOAC: 0.36°/6.36 cm/2.96 ms and for MOISST: 2.24°/56.34 cm/27.07 ms. Detailed quantitative results by sequence are given in the supplementary materials.
Table 2: LiDAR/Camera calibration results on KITTI-360.
Method Rotation (°) Translation (cm)
SOAC $\mathbf{0.33}$ $\mathbf{7.75}$
LCCNet $1.92$ $95.8$
@@
For the task of LiDAR/Camera calibration, we compare our method against LCCNet [20]. The provided weights were pretrained on the KITTI odometry dataset [8]. We predict the calibration between the Front-Left camera and the LiDAR of KITTI-360, as it is a setup very similar to KITTI odometry. We can see in Table 2 that the performance of LCCNet, especially on the translation, is very poor compared to SOAC (results per sequence are provided in the supplementary materials). As it is a supervised method, we observe that it is setup-specific, and a slight change in the LiDAR/Camera configuration greatly reduces the performance. This was also highlighted in the work by Fu et al. [7] when using the Front-Right camera for calibration on the KITTI odometry dataset.
Refer to caption Refer to caption Refer to caption
Figure 6: Results on nuScenes with 5 cameras for SOAC and SOAC w/o NeRF delaying as box plots with log scale, the red lines show the initial error (Best viewed in color).
@@
Figure 7: Results on Pandaset for SOAC, MOISST and SOAC w/o semantic filtering as box plots with log scale, the red lines show the initial error (Best viewed in color).
Calibration in dynamic environment.
For the evaluation on dynamic environment, we use 3 Pandaset sequences with the presence of dynamic elements such as cars or pedestrians. When calibrating on dynamic scenes, the moving elements are not handled by the NeRF model. However, a simple and efficient way of removing these elements is to filter the dynamic classes with semantic segmentation. We lose some useful information for calibration, for instance stationed cars, but if the rest of the scene is sufficiently informational, proper calibration can be obtained. We apply a similar setup to the one for KITTI-360 and nuScenes, except that we remove the temporal calibration and the initial time offset (cf. Section 4.3 on Time-space compensation). We downscale the image by a factor of 4 for SOAC and 2 for MOISST. We use semantic segmentation computed by Mask2Former [4] to remove all classes that can be considered dynamic. We show the results of SOAC vs. MOISST and SOAC without semantic filtering in Figure 7. By using the semantic filtering, we greatly improve the calibration performance on the LiDAR with a median error for SOAC of 0.41°/7.79 cm compared to MOISST: 2.36°/79.17 cm. We can also see that SOAC performs much better than MOISST on the overall calibration of all the sensors with error an of 0.42°/11.18 cm versus 2.18°/44.73 cm.
@@
To evaluate SOAC performances with a nearly complete 360° camera rig, we add two additional side cameras on the nuScenes sequences. We run both with and without the NeRFs delaying scheduling as explained in section 3.6. In Figure 6 we can see the impact of not delaying the NeRFs, as the accuracy and stability of the calibration plummet.
@@
Figure 9: Qualitative LiDAR/front camera calibration results on KITTI-360 dataset.
@@
Refer to caption Refer to caption Refer to caption
SOAC
@@
Figure 10: Qualitative LiDAR/Camera reprojection results on nuScenes dataset.
We show the reprojection of the LiDAR on the images using the calibration obtained from different methods. On KITTI-360 (cf. Figure 9) we can see that LCCNet does not provide a satisfying result and that SOAC is able to provide a visually comparable alignment to the ground truth calibration. On nuScenes (cf. Figure LABEL:fig:reproj_nuScenes), the calibration from SOAC provides a better alignment than MOISST, assessing the quantitative results of Figure 5. More qualitative results are given in the supplementary materials, along with ablation studies on visibility grids (cf. Section 3.5) and correction bounding (cf. Section 3.6).
@@ 4.3 Limitations
Table 3: Illustration of SOAC space-time compensation on a sequence from KITTI-360 – Mean absolute poses of sensors are correct whereas the spatio-temporal calibration computed by the method is erroneous.
@@
Time-space compensation.
When simultaneously calibrating spatially and temporally, there are cases where the disentanglement is impossible. In a sequence where the vehicle is driving in a straight line at a constant speed, there is an infinite number of solutions that can provide the correct poses. In Table 3, we show the calibration results on a straight line with constant speed from KITTI-360. We can see the pose error is fairly low, but the extrinsic and temporal calibration is incorrect. This means that there is a need to select scenes with speed variation in order to reduce to a single possible solution. As most Pandaset sequences are in a straight line at a constant speed, we decided to not do temporal calibration on them.
Scene structure.
When the scenes are more open and/or larger, the projected rays will have to travel a longer distance before reaching the scene’s structures. Considering LiDAR to camera calibration, the rotation error has a linearly increasing impact according to the ray distance when reprojected to the camera frame, while the translation error’s impact is independent from the ray distance. Thus, we tend to lose precision on the translation as the ray gets longer. When analyzing the LiDAR rays length distribution of the datasets in Figure 8, we observe that the LiDAR rays on Pandaset are longer, meaning that the scenes are larger and open, and the structures are farther than on KITTI-360 and nuScenes. This explains most likely the decrease in calibration performances for the LiDAR extrinsic translation parameters on Pandaset (median error of 18.1 cm) compared to KITTI-360 (median error of 8.2 cm) or nuScenes (median error of 5.1 cm).
Training time.
As we train one NeRF per camera, and register all the other sensors on each NeRF, the training time increases exponentially with the number of cameras. For instance, on nuScenes one epoch takes approximately 1 minute 45 seconds for 3 cameras and 8 minutes for 5 cameras using the same GPU. This reduces the scalability of our method, but this phenomenon is mitigated by the fact that we use much smaller images than MOISST to reach better performance (see supplementary materials for details about the total training time of the methods).
@@ A.1 Datasets
KITTI-360 [18]: Sequences are selected and cropped by considering vehicle speed variations to remove time-space compensation issues as described in main article Section 4.3. Once sequences are cropped, one out of two frames are kept for all sensors to obtain a total of 40 frames per sequence. This decision was made to match the same length as the NVS benchmark sequences present on the dataset. The details from each sequence are summarized in Table 4.
Table 4: Selected frames for each KITTI-360 sequence.
@@
nuScenes [3]: Since nuScnes poses are provided only in $\mathrm{SE(2)}$, they cannot be used directly for our method. Instead, we use KISS-ICP [37] to get a good estimate of the LiDAR poses. Extrinsic calibration provided by the dataset is then used to obtain the poses for all cameras. We select the sequences 916, 410 and 417 for our experiments, as they are more suitable for the calibration (closer structures, more speed variation). All LiDAR scans are used during calibration as the LiDAR is sparser than the one in KITTI-360, while one out of two images is subsampled to reduce training time.
Pandaset [42]: Since extrinsic parameters are not provided by the dataset, they are estimated using the global poses of all sensors at several frames by calculating the transformation between the frames with the same timestamp from each sensor. Sequences 33, 40 and 53 are used for our experiments as they have more close structures. We apply the same subsampling strategy as for nuScenes.
@@ A.2 Architecture and Losses
Figure 11: Ablation results on KITTI-360 sequence 4: for SOAC, SOAC w/o Sigmoid and SOAC w/o visibility grid as box plots with log scale, the red lines show the initial error (Best viewed in color).
@@ A.3 Hyperparameters
Table 6: Dataset specific SOAC hyperparameters used for the training (in epochs).
Hyperparameter KITTI-360 nuScenes Pandaset
Diagonal cams NeRF delay - 1 3
Side cams NeRF delay 1 9 -
LiDAR delay 6 5 8
In Table 5 are indicated the hyperparameters used for the training of SOAC, and in Table 6 the hyperparameters specific for each dataset. The dataset-specific parameters largely depend on the overlap between the sensors, and the frame number/size. Basically, larger overlaps and larger quantities of data reduce the number of delay epochs.
The number of epochs for training MOISST is reduced to 20, as improvement was not observed with more. The spatial and temporal optimization learning rate is fine-tuned to 5e-4.
@@ Appendix B Additional ablations
The addition of the sigmoid for bounding the translation and temporal corrections allows better stability and robustness as shown in Figure 11 on which a huge decrease in calibration accuracy can be noticed when removing the sigmoid.
@@
Removing the visibility grids deteriorates the performance of the LiDAR calibration rotation and translation as shown in Figure 11.
Appendix C Training time
Table 7: Training time comparison on different sequences.
Dataset MOISST SOAC
KITTI-360 $\sim$ 2 h 30 min $\sim$ 1 h 30 min
Nuscenes 3 cams $\sim$ 2 h 30 min $\sim$ 1 h
Nuscenes 5 cams $\sim$ 4 h 30 min $\sim$ 2 h 30 min
Pandaset $\sim$ 1 h 30 min $\sim$ 1 h 20 min
In Table 7 are indicated the training time for SOAC and MOISST with a high-tier GPU on the sequences. The indicated results are with the downscaled images as described in the paper. SOAC is able to provide better calibration than MOISST with shorter training time, even if multiple NeRFs are used, as it can use much smaller images.
Appendix D Quantitative results
Specific box plot results are provided for each sequence. The results for KITTI-360 are in Figure 12, the results for Nuscenes in Figure 13, and the results for Pandaset in Figure 14.
@@ Appendix D Quantitative results
Figure 12: Results for KITTI-360 per sequence for SOAC and MOISST as box plots with log scale. The red lines show the initial error (Best viewed in color).
@@
Figure 13: Results for nuScenes per sequence for SOAC and MOISST as box plots with log scale. The red lines show the initial error (Best viewed in color).
Sequence 33 Sequence 40 Sequence 53
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 14: Results for Pandaset per sequence for SOAC and MOISST as box plots with log scale. The red lines show the initial error (Best viewed in color).

### 27 Mar 2024 v3

@@ 3.6 Optimization Details
Overall, the training process can be summarized as follows: during each training step, a mini-batch of rays is first used in the scene representation training step (Sec. 1(a)). Rays of each camera train their specific NeRF and fill the respective visibility grids (Sec. 3.5). The LiDAR rays train all the NeRFs and are used to optimize the LiDAR calibration parameters after being filtered by the visibility grids. In a subsequent step, the same mini-batch is passed to the extrinsic and temporal optimization. Rays are filtered through the visibility grids before being fed to the NeRFs as explained in Sec. 3.3. Calibration losses (Eq. 8) are computed and the gradient is backpropagated to optimize the calibration parameters. Once this is done, we continue the training with the next mini-batch.
@@
Figure 5: Results for SOAC and MOISST [12] as box plots with log scale on KITTI-360 [18] and Nuscenes [3] sequences. The red lines show the initial error (best viewed in color).
@@ 4.1 Setup
We perform experiments on three popular autonomous driving datasets: KITTI-360 [18], nuScenes [3] and Pandaset [44]. For KITTI-360, we use the two front cameras, the two side cameras and the Velodyne LiDAR for our experiments. For nuScenes and Pandaset, we use the front camera, the two front diagonal cameras, and the LiDAR. Undistorted LiDAR scans are considered for all datasets. We assign the front-left camera of KITTI-360, and the front cameras of nuScenes and Pandaset to be the reference sensor. More details on selected sequences and dataset parameters are provided in the supplementary materials.
Refer to caption Refer to caption Refer to caption
Figure 6: Results on nuScenes [3] with 5 cameras for SOAC and SOAC w/o NeRF delaying as box plots with log scale, the red lines show the initial error (best viewed in color).
@@
We select MOISST [12] as our baseline, as it also aims to solve targetless, multi-modal, and spatiotemporal calibration. We refer to the supplementary for details about the re-implementation of the method. For the LiDAR/Camera calibration task, we compare against LCCNet [20] by using the code and the pre-trained weights from the official repository [^1] https://github.com/IIPCVLAB/LCCNet, and with Pandey et al. [29] using the official implementation provided by authors [^2] https://robots.engin.umich.edu/SoftwareData/InfoExtrinsicCalib. For SOAC, We replicate MOISST NeRF architecture and apply the same supervision and regularization losses. We refer to the supplementary for more implementation details.
@@ 4.2 Results
We run both SOAC and MOISST on 4 KITTI-360 sequences and 3 nuScenes sequences. For SOAC, KITTI-360 images are downscaled by 4, while a downscale factor of 6 is applied for nuScenes. For MOISST, we do not downscale the KITTI-360 images and apply a downscale factor of 2 for nuScenes as we found that the method performs better with high-resolution images. Each test is run with an initial noise of 50 cm translation error and 5° rotation error on each axis, as well as 100 ms of time offset. We use 10 different seeds to randomly sign the error noises applied and compute the statistics over these 10 runs. Following common practices [37, 33], we show results on Fig. 5 by employing box plots [^3] The boxes show the first quartile $Q_{1}$, median, third quartile $Q_{3}$. The whiskers use 1.5 IQR (Interquartile range) above and below the box and stop at a value within the results. As can be seen, SOAC achieves better calibration results on KITTI-360 with an overall error (average over median for each sensor) of 0.21°, 5.24 cm and 3.95 ms for rotation, translation and time offset, respectively. In contrast, MOISST obtains errors about 10 times higher (i.e. 2.24°, 56.34 cm and 27.07 ms) for the same setup. Detailed quantitative results by sequence are given in the supplementary materials.
KITTI-360 [18] Pandaset [44]
Rotation (°) Translation (cm) Rotation (°) Translation (cm)
Pandey et al. [29] $11.8\pm 5.4$ $143\pm 109$ $15.4\pm 0.8$ $139\pm 17.5$
LCCNet [20] $1.9\pm 0.1$ $95.8\pm 7.7$ $14.3\pm 3.4$ $370\pm 11.6$
MOISST [12] $\mathbf{0.2\pm 0.1}$ $10.0\pm 9.8$ $2.8\pm 2.3$ $56.4\pm 17.2$
SOAC (ours) $0.3\pm 0.2$ $\mathbf{7.8\pm 3.5}$ $\mathbf{1.3\pm 0.8}$ $\mathbf{29.4\pm 13.6}$
Table 2: LiDAR/Camera calibration results.
@@
For the task of LiDAR/Camera calibration, the same initial rotation and translation error setup from previous experiments is applied, but without considering any temporal error. We compare our method against LCCNet [20] and Pandey et al. [29]. The provided weights for LCCNet were pre-trained on the KITTI odometry dataset [8]. For KITTI-360, We predict the calibration between the front-left camera and the LiDAR. For Pandaset, we predict the calibration between the front camera and the 360° LiDAR. Results are shown in Tab. 2. The performance of LCCNet, is very poor in comparison to SOAC, especially for the translation (results per sequence are provided in the supplementary). As LLCNet is a supervised method, we observe that it is setup-specific, and a slight change in the LiDAR/Camera configuration greatly reduces the performance. This was also highlighted by Fu et al. [7] when using the front-right camera for calibration on the KITTI odometry dataset. For Pandey et al. [29], we were unable to obtain convincing calibration results on the sequences. We argue that feature-based targetless methods are not designed for “in-the-wild” calibration, and sequences need to be acquired in a specific manner to obtain proper results (i.e. indoor, structured environment, dense LiDAR).
@@
Figure 7: Results on Pandaset [44] for SOAC, MOISST [12] and SOAC w/o semantic filtering as box plots with log scale, the red lines show the initial error (best viewed in color).
Calibration in dynamic environments.
For the evaluation in dynamic environments, we select 3 Pandaset sequences with the presence of dynamic elements (e.g. cars, pedestrians). When calibrating on dynamic scenes, the moving elements are not handled by the NeRF model. Therefore, a simple and efficient way of removing these elements is to filter the dynamic classes with semantic segmentation. This results in losing some useful information for calibration (i.e. parked vehicles). Nevertheless, if the rest of the scene provides sufficient overlap, proper calibration can be obtained. We apply an analogous setup to KITTI-360 and nuScenes, except for the removal of temporal calibration and the initial time offset (cf. Sec. 4.3 on Time-space compensation). We downscale the image by a factor of 4 for SOAC and 2 for MOISST. We use semantic segmentation computed by Mask2Former [4] to remove all classes that can be considered dynamic for both methods and test SOAC with and w/out semantic filtering. Results are shown in Fig. 7. It can be observed that by applying semantic filtering, calibration performance on SOAC can be greatly improved on the LiDAR with a median error of 0.41° and 7.79 cm on rotation and translation, respectively, in comparison to results w/out filtering (2.36° / 79.17 cm). It can be also seen that SOAC performs much better than MOISST on the overall calibration of all the sensors with a mean error an of 0.42° / 11.18 cm vs. 2.18° / 44.73 cm for MOISST.
@@
To evaluate SOAC performances with a nearly complete 360° camera rig, we add two additional side cameras on the nuScenes sequences. We run both with and without the NeRFs delaying scheduling as explained in Sec. 3.6. In Fig. 6 we can see the impact of not delaying the NeRFs, as the accuracy and stability of the calibration plummet.
@@
SOAC (ours)
Refer to caption Refer to caption
Figure 9: Qualitative LiDAR/front camera calibration results on KITTI-360 [18] dataset.
@@
SOAC (ours)
Refer to caption Refer to caption Refer to caption
Figure 10: Qualitative LiDAR/Camera reprojection results on nuScenes [3] dataset.
Qualitative results.
We show the reprojection of the LiDAR on the images using the calibration obtained from different methods. On KITTI-360 (cf. Fig. 9) we can see that LCCNet does not provide a satisfying result and that SOAC is able to provide a visually comparable alignment to the ground truth calibration. On nuScenes (cf. Fig. 10), the calibration from SOAC provides a better alignment than MOISST, assessing the quantitative results of Fig. 5. More qualitative results are given in the supplementary, along with ablation studies on visibility grids (cf. Sec. 3.5) and correction bounding (cf. Sec. 3.6).
@@ 4.3 Limitations
Table 3: SOAC space-time compensation on a sequence from KITTI-360 [18]. Mean absolute poses of sensors are correct whereas the spatio-temporal calibration computed by the method is erroneous.
Time-space compensation. When simultaneously calibrating spatially and temporally, there are cases where the disentanglement is impossible. In a sequence where the vehicle is driving in a straight line at a constant speed, there is an infinite number of solutions that can provide the correct poses. In Tab. 3, we show the calibration results on a straight line with constant speed from KITTI-360. We can see the pose error is fairly low, but the extrinsic and temporal calibration is incorrect. This means that there is a need to select scenes with speed variation in order to reduce to a single possible solution. As most Pandaset sequences are in a straight line at a constant speed, we decided to not do temporal calibration on them.
Scene structure. When the scenes are more open and/or larger, the projected rays will have to travel a longer distance before reaching the scene’s structures. Considering LiDAR to camera calibration, the rotation error has a linearly increasing impact according to the ray distance when reprojected to the camera frame, while the translation error’s impact is independent of the ray distance. Thus, we tend to lose precision on the translation as the ray gets longer. When analyzing the LiDAR rays length distribution of the datasets in Fig. 8, we observe that the LiDAR rays on Pandaset are longer, meaning that the scenes are larger and open, and the structures are farther than on KITTI-360 and nuScenes. This explains most likely the decrease in calibration performances for the LiDAR extrinsic translation parameters on Pandaset (median error of 18.1 cm) compared to KITTI-360 (median error of 8.2 cm) or nuScenes (median error of 5.1 cm).
Training time. As we train one NeRF per camera, and register all the other sensors on each NeRF, the training time increases exponentially with the number of cameras. For instance, on nuScenes one epoch takes approximately 1 minute 45 seconds for 3 cameras and 8 minutes for 5 cameras using the same GPU. This reduces the scalability of our method, but this phenomenon is mitigated by the fact that we use much smaller images than MOISST to reach better performance. We refer to the supplementary for more in-depth details on the efficiency of our method wrt. image size compared to MOISST.
@@ A.1 Datasets
KITTI-360 [18]:
Sequences are selected and cropped by considering vehicle speed variations to remove time-space compensation issues as described in main article Sec. 4.3. Once sequences are cropped, one out of two frames are kept for all sensors to obtain a total of 40 frames per sequence. This decision was made to match the same length as the NVS benchmark sequences present on the dataset. The details from each sequence are summarized in Tab. 4.
@@
Table 4: Selected frames for each KITTI-360 [18] sequence.
nuScenes [3]:
Since nuScnes poses are provided only in $\mathrm{SE(2)}$, they cannot be used directly for our method. Instead, we use KISS-ICP [39] to get a good estimate of the LiDAR poses. Extrinsic calibration provided by the dataset is then used to obtain the poses for all cameras. We select the sequences 916, 410 and 417 for our experiments, as they are more suitable for the calibration (closer structures, more speed variation). All LiDAR scans are used during calibration as the LiDAR is sparser than the one in KITTI-360, while one out of two images is subsampled to reduce training time.
Pandaset [44]:
Since extrinsic parameters are not provided by the dataset, they are estimated using the global poses of all sensors at several frames by calculating the transformation between the frames with the same timestamp from each sensor. Sequences 33, 40 and 53 are used for our experiments as they have more close structures. We apply the same subsampling strategy as for nuScenes.
@@ A.2 Architecture and Losses
Figure 11: Ablation results on KITTI-360 [18] sequence 4: for SOAC, SOAC w/o Sigmoid and SOAC w/o visibility grid as box plots with log scale, the red lines show the initial error (Best viewed in color).
@@ A.3 Hyperparameters
Table 5: SOAC hyperparameters used for the training.
Sensor KITTI-360 [18] nuScenes [3] Pandaset [44]
Diagonal cams - 1 3
Side cams 1 9 -
LiDAR 6 5 8
Table 6: SOAC hyperparameters used for the training.
In Tab. 6 are indicated the hyperparameters used for the training of SOAC, and in Tab. LABEL:tab:dataset_hyperparameters are the NeRF delaying epochs depending on the dataset. Delaying the NeRF proves advantageous in scenarios characterized by a multitude of sensors, some of which exhibit minimal overlap with the reference sensor throughout the sequence. This approach facilitates the accurate propagation of calibration information during training from sensors presenting significant overlap with the reference sensor to those with lesser or no overlap at all. Basically, larger overlaps and larger quantities of data reduce the number of necessary delay epochs. The number of epochs for training MOISST is reduced to 20, as improvement was not observed with more. The spatial and temporal optimization learning rate is fine-tuned to 5e-4.
@@ Appendix B Additional ablations
The addition of the sigmoid for bounding the translation and temporal corrections allows better stability and robustness as shown in Fig. 11 on which a huge decrease in calibration accuracy can be noticed when removing the sigmoid.
@@
Removing the visibility grids deteriorates the performance of the LiDAR calibration rotation and translation as shown in Fig. 11.
@@
Figure 12: Results for KITTI-360 [18] per sequence for SOAC and MOISST [12] as box plots with log scale. The red lines show the initial error (best viewed in color).
Sequence 33 Sequence 40 Sequence 53
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 13: Results for Pandaset [29] per sequence for SOAC and MOISST [12] as box plots with log scale. The red lines show the initial error (best viewed in color).
@@
Figure 14: Results for nuScenes [3] per sequence for SOAC and MOISST [12] as box plots with log scale. The red lines show the initial error (best viewed in color).
Appendix C Training time
Dataset MOISST [12] SOAC
KITTI-360 [18] $\sim$ 2 h 30 min $\sim$ 1 h 30 min
Nuscenes [3] (3 cams) $\sim$ 2 h 30 min $\sim$ 1 h
Nuscenes [3] (5 cams) $\sim$ 4 h 30 min $\sim$ 2 h 30 min
Pandaset [44] $\sim$ 1 h 30 min $\sim$ 1 h 20 min
Table 7: Training time comparison on different sequences.
  Downscale Calibration error(°/cm/ms) Training time (min)
factor SOAC MOISST [12] SOAC MOISST [12]
1 0.2/4.6/3.9 0.1/5.3/1.3 605 163
2 0.3/4.6/2.5 0.3/24.1/5.3 181 42
4 0.2/4.6/1.7 1.1/41.4/13.1 85 17
8 0.4/12.3/8.2 2.6/56.6/28.6 53 12
Table 8: Training time and calibration accuracy for varying downscale factor on KITTI-360 [18] sequence 1 seed 0
We report the mean training times with both SOAC and MOISST for the sequences from each tested dataset in Tab. 7. For all the experiments, we used a GPU of similar performance to an RTX 3090. The shown results are with the downscaled images as described in the paper. SOAC is able to provide better calibration than MOISST with shorter training time, even if multiple NeRFs are used, as it can use much smaller images. To measure the impact of the image downscale factor in relation to each method’s training time, we train both methods at different downscale factors and report results on Tab. 8. As it can be observed, MOISST accuracy is considerably harmed by using lower-resolution images in comparison to SOAC. Furthermore, SOAC achieves high accuracy even with large downscale factors on the images (i.e. downscaling the image resolution by 4 shows no drop in accuracy for SOAC while being 8 times faster. In comparison, MOISST presents a severe drop in performance when downscaling). This enables SOAC to achieve more efficient training times given its ability to exploit lower-resolution images.
Appendix D Quantitative results
Specific box plot results are provided for each sequence. The results for KITTI-360 are in Fig. 12, the results for Nuscenes in Fig. 14, and the results for Pandaset in Fig. 13.